OpenStudy (anonymous):

must simplify [3cos(x) - 12sin^2(x)cos(x)] into 3cos(3x), help please

OpenStudy (anonymous):

lets see other reply s...

OpenStudy (anonymous):

hmm i can get so many variations but no idea how to get 3cos(3x)

OpenStudy (anonymous):

oh yes it is !

OpenStudy (anonymous):

can yu please show me how!! i know the rule cos (x+y) or something is in it i think but im just so lost

OpenStudy (callisto):

Wow! \[3cosx - 12sin^2 x cosx\]\[= 3 ( cosx -4sin^2 xcosx)\]\[=3[cosx - 2(2sinxcosx)sinx]\]\[=3(cosx-2sin2xsinx)\]\[=3[cosx-2(-\frac{1}{2}cos(2x+x)-cos(2x-x))]\]simplify ... Amazing :D

OpenStudy (anonymous):

i got logged out, sorry

OpenStudy (anonymous):

ok first off we need an expression for \(\cos(3x)\) in terms of \(\cos(x)\) then we can write the above in terms of \(\cos(x)\) and see if they match up that will be the strategy i think

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] and so \[\cos(3x)=\cos(2x+x)=\] actually @Callisto way a lot shorter, nvm

OpenStudy (anonymous):

but if you can start with the fact that \[\cos(3x)=4\cos^3(x)-3\cos(x)\] then it is some simple algebra if you have to derive that formula, it takes a bunch of steps

OpenStudy (anonymous):

nice work @Callisto

OpenStudy (callisto):

It's not shorter... just different approach... \[3cos3x\]\[=3cos(2x+x)\]\[=3(cos2xcosx-sin2xsinx)\]\[=3(1-2sin^2x)cosx - 3(2sinxcosx)(sinx)\]\[=3cosx-6sin^2xcosx-6sin^2xcosx\]\[=...\]Amazing :) I should have thought from the result :|

OpenStudy (anonymous):

thanks guys!